# Towards Context-Aware Domain Generalization: Understanding the Benefits and Limits of Marginal Transfer Learning

### v2

@@ 5 Conclusions
In this work, we aimed to advance the theoretical understanding of marginal transfer learning in domain generalization. Thus, we formulated criteria that are necessary for the approach to yield benefits and are also easy to verify empirically. Moreover, we pinpoint the source component shift as a scenario where context-aware models can offer advantages, enabling the identification of favorable scenarios. Additionally, we illustrate the capability to identify novel environments, thereby allowing for the detection of potential failure cases. This, in turn, also supports model selection by choosing the most predictive on ID data and the most robust model on OOD data.
Acknowledgments
JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation. We thank Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
Appendix A Algorithm
@@
for $i=1,\dots,k$ do
       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
       for $j=1,\dots,m$ do
             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
       end for
      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
@@
A.1 Model Overview
For a detailed overview of the models used to verify the criteria, see Table 3.
Model Symbol Description Purpose
Context-aware model (ours) $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts Y from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Improve predictions
Baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$ Predicts Y from $\mathbf{X}$ alone Verify improvement
Environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ Predicts Y from $\mathbf{X}$ and $E$ Verify Criterion 2.3
Contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $E$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Verify Criterion 2.2
Baseline environment model $f^{E\mid\mathbf{X}}$ Predicts $E$ from $\mathbf{X}$ alone Verify Criterion 2.2.
Table 3: The five different model types used to evaluate our approach and verify the theoretical criteria for improvement.
@@ Appendix B Additional Experiment: ProDAS
We utilize the ProDAS library [Müller et al., 2023] to generate high-dimensional image data that meets our dataset requirements. The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size. Additionally, the background varies in color and texture, resulting in a complex scenario. For examples see Figure 9. We consider the task of predicting the object size. Difficulties arise due to the presence of distinct environments with varying characteristics. Specifically, depending on the environment, a constant is added to the observed object size to get the actual target variable that we aim to predict:
@@
Here, $j\in\{1,2,3,4\}$ denotes the environment, while $Y_{\text{gt}}$ represents the ground truth (or factual) size, obtained as a sum of the observed size $Y_{\text{observed}}$ (relative to the image frame) and a constant depending on $j$. The background color follows a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{j};\boldsymbol{\Sigma})$ where the mean depends on the environment in the following way: $\boldsymbol{\mu}_{j}=\boldsymbol{\mu}_{0}+j\cdot\text{const}_{2}$. Here we assign a small value to $\text{const}_{2}$ to enforce the background distributions to overlap between different environments. Specifically, this construction implies that the relation between input $\mathbf{X}$ and target $Y$ differs across environments. This corresponds to Criterion 2.3. Notably, inferring the originating environment from a single sample is unattainable due to overlapping background distributions (corresponding to Criterion 2.2). Samples of different environments are shown in Appendix F. This example could be inspired by microscopy data where different microscopes correspond to distinct environments, each exhibiting its own characteristics. During training, we assume to have access to the ground truth value $Y_{\text{gt}}$.
@@ Appendix C Theory
The proof can be found in subsection C.3.
@@ C.2 Insufficiency of Criteria 2 and 3 for Criterion 1
Regarding the mechanism relating inputs to outputs, we assume that on $[a,\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ differs, e.g., two constant functions with distinct values. We further assume that on $(\frac{a+b}{2},b+\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ does not vary with the environment, e.g., is constant. This aligns with Criterion 2.3: if we know the environment, we can improve the prediction, specifically on $[a,\frac{a+b}{2}]$.
However, Criterion 2.1 is not satisfiable. The set input allows us to distinguish environment 3 (i.e. the one with support $\mathcal{U}[a+\frac{a+b}{2},b+\frac{a+b}{2}]$) from the other ones. Yet, we cannot distinguish between environment 1 and environment 2. Since the relation between $X$ and output $Y$ differs only in the supports of environment 1 and environment 2 (specifically, it differs in $\mathcal{U}[a,\frac{a+b}{2}]$), the set input cannot provide additional information about the output $Y$ compared to the single input $X$, i.e. it holds $Y\perp\mathbf{S}^{(n)}\mid X$,
C.3 Proof of Theorem 2.1
@@
=$\overset{(\star)}{=}\left(h(A)+h(B)-h(A\mid A+B)\right)-\left(h(A\mid C)+h(B\mid C)-h(A\mid A+B,C)\right)$
@@
()($\star$)( ⋆ ) follows with the chain rule for entropy
@@
[…]
subsection C.3 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
@@
In the following, we discuss the assumptions in (c) and (d). In our experiments, we observed that in most datasets a relatively small sample size suffices to infer the environment label with approximately 100% accuracy (see Table 6). Therefore, the assumption that there exists a function $g(\mathbf{S}^{(n)})=E$ seems justified if $n$ is sufficiently large. To generalize the assumption where the environment label is not fully inferable, we have to make assumptions. For one, we require $\mathbf{S}^{(n)}\mid Z\mid\mathbf{X}$. This can be interpreted as “increasing the set size does not improve the prediction of $E$”. Also $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$ can be interpreted similarly: increasing the set size and considering the ground truth label/value does not enhance the predictability of $E$. Both assumptions should hold approximately if $n$ is large enough. With the assumption $I(\mathbf{Y};E\mid\mathbf{X})>I(Z;\perp\mathbf{Y}\mid\mathbf{X})$ we assume that the noise $Z$ is less predictive of $Y$ compared to $E$ if $\mathbf{X}$ is given. This can be roughly interpreted as the noise does not prove useful for predicting $Y$ from $\mathbf{X}$ compared to the ground truth environment label.
@@ Appendix D Experiments: General Remarks
$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X}})}$ (22)
@@
$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}$ (23)
@@ E.1 Data Generation
Simpson’s Paradox [Peters et al., 2017, von Kügelgen et al., 2021] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined. There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [Bickel et al., 1975] or an evaluation of the efficacy of different treatments for kidney stones [Charig et al., 1986].
To replicate this, we create a dataset inspired by an illustration of Simpson’s Paradox on Wikipedia sim . The dataset consists of a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target. Unless otherwise specified, we generate the data by taking an equal number of samples from each mixture component, defining the environment as a one-hot vector over the mixture components.
@@
n_domains $5$ number of mixture components
@@
The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [minutephysics and Reich, 2017].
@@ E.3 Non-Linear Models
In the experiments in subsection 4.2, we considered linear models for our model and the baseline. In the following, we show results for the non-linear model class in Figure 7. We compare predictions of a baseline model and our model on all environments in Figure 8. We see that the extrapolation task fails in some cases as in environment 1. This is due to the mismatch of the considered model class and ground truth model.
@@ Appendix G Experiment 2: Details
To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data. Therefore, we follow the approach proposed in subsection 2.5. We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$. Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension. To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see subsection 2.5 for details).
@@ Appendix H Experiment 3 and 4: Details
For the BikeSharing dataset we consider a simple feed-forward neural network in all models. For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized. Here, we employ the Clip features proposed in Radford et al. [2021]. The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions. Given that Clip features considerably simplify the task, we performed a minimal hyper-parameter search and ensured that the inference model had a similar number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
Table 5: Experiment 4. Performance comparison between our model and the baseline, broken down by target domain. We compare their performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case. We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
@@
Dataset / Set Size ProDAS / $128$ OfficeHome / $4$ PACS / $4$
@@ Appendix I Comparison of Permutation-Invariant Architectures
Therefore, we construct the toy dataset as described in subsection E.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
@@
Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [Agarap, 2019] activation. For the set-based methods, this is followed by the respective pooling. We choose $8$ heads for the attention-based model.
Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [Zhang and Sabuncu, 2018].
@@
This dataset, taken from the UCI machine learning repository [Fanaee-T, 2013], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.

### v3

@@ 5 Conclusions
In this work, we aimed to advance the theoretical understanding of marginal transfer learning in domain generalization. Accordingly, we formalized criteria that are necessary for context-aware models to yield benefits and are also verifiable in practice. Moreover, we pinpointed the source component shift as a scenario where context-aware models can offer advantages, enabling the identification of favorable settings and the identification of potential failure cases. The latter allows us to perform real-time model selection between the best performing model on ID data and the most robust (i.e., domain-invariant) model on OOD data. Future research should investigate generalization bounds, the learner’s behavior in finite-data regimes, and the generalization behavior of the learner as the number of training domains increases (i.e., the domain efficieincy).
Acknowledgments and Disclosure of Funding
This work is partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project 528702768. JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation. We thank Paul Christian Bürkner, Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
Appendix A Pseudocode
@@
1 for $i=1,\dots,k$ do
2   Sample mini-batch $\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
3 for $j=1,\dots,m$ do
4      Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
5      Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
6   end for
7  Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
@@ Appendix B Additional Experiment: ProDAS
We utilize the ProDAS library [58] to generate high-dimensional image data that meets our dataset requirements. The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size. Additionally, the background varies in color and texture, resulting in a complex scenario. For examples see Figure 10. We consider the task of predicting the object size. Difficulties arise due to the presence of distinct environments with varying characteristics. Specifically, depending on the environment, a constant is added to the observed object size to get the actual target variable that we aim to predict:
@@
Here, $j\in\{1,2,3,4\}$ denotes the environment, while $Y_{\text{gt}}$ represents the ground truth (or factual) size, obtained as a sum of the observed size $Y_{\text{observed}}$ (relative to the image frame) and a constant depending on $j$.
The background color follows a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{j};\boldsymbol{\Sigma})$ where the mean depends on the environment in the following way: $\boldsymbol{\mu}_{j}=\boldsymbol{\mu}_{0}+j\cdot\text{const}_{2}$. Here we assign a small value to $\text{const}_{2}$ to enforce the background distributions to overlap between different environments. Specifically, this construction implies that the relation between input $\mathbf{X}$ and target $Y$ differs across environments. This corresponds to Criterion 2.3. Notably, inferring the originating environment from a single sample is unattainable due to overlapping background distributions (corresponding to Criterion 2.2). Samples of different environments are shown in Appendix F. This example could be inspired by microscopy data where different microscopes correspond to distinct environments, each exhibiting its own characteristics. During training, we assume to have access to the ground truth value $Y_{\text{gt}}$.
@@ Appendix C Theory
The proof can be found in Section C.3.
@@ C.2 Insufficiency of Criteria 2 and 3 for Criterion 1
Regarding the mechanism relating inputs to outputs, we assume that on $a,\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ differs, e.g., two constant functions with distinct values. We further assume that on $(\frac{a+b}{2},b+\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ does not vary with the environment, e.g., is constant. This aligns with Criterion 2.3: if we know the environment, we can improve the prediction, specifically on $a,\frac{a+b}{2}]$.
However, Criterion 2.1 is not satisfiable. The set input allows us to distinguish environment 3 (i.e. the one with support $\mathcal{U}[a+\frac{a+b}{2},b+\frac{a+b}{2}]$) from the other ones. Yet, we cannot distinguish between environment 1 and environment 2. Since the relation between $X$ and output $Y$ differs only in the supports of environment 1 and environment 2 (specifically, it differs in $\mathcal{U}[a,\frac{a+b}{2}]$), the set input cannot provide additional information about the output $Y$ compared to the single input $X$, i.e. it holds $Y\perp\mathbf{S}^{(n)}\mid X$.
It is also worth noting that Criterion 2.3 might be achievable while Criterion 2.2 is unattainable and vice versa. For instance, when we can infer the originating environment from one sample (Criterion 2.2 is not attainable), the relation between $\mathbf{X}$ and $\mathbf{Y}$ might still vary with the environment (Criterion 2.3 is achievable).
C.3 Illustration and Proof of Theorem 2.1
Figure 6: Illustration of Theorem 2.1. The first row depicts (a), the second row (b) and the third row (c). The pink framed plots show the conditional distributions along the pink marker as shown on the right.
@@
$\overset{(\star)}{=}\left(h(A)+h(B)-h(A\mid A+B)\right)-\left(h(A\mid C)+h(B\mid C)-h(A\mid A+B,C)\right)$
@@
$(\star)$ follows with the chain rule for entropy
@@
$(\star\star)$ follows since $h(A\mid A+B,C)\leq h(A\mid A+B)$.
Section C.3 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
@@
In the following, we discuss the assumptions in (c) and (d). In our experiments, we observed that in most datasets a relatively small sample size suffices to infer the environment label with approximately 100% accuracy (see Table 6). Therefore, the assumption that there exists a function $g(\mathbf{S}^{(n)})=E$ seems justified if $n$ is sufficiently large. To generalize the assumption where the environment label is not fully inferable, we have to make assumptions. For one, we require $\mathbf{S}^{(n)}\mid Z\mid\mathbf{X}$. This can be interpreted as “increasing the set size does not improve the prediction of EE”. Also $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$ can be interpreted similarly: increasing the set size and considering the ground truth label/value does not enhance the predictability of $E$. Both assumptions should hold approximately if $n$ is large enough. With the assumption $I(\mathbf{Y};E\mid\mathbf{X})>I(Z;\perp\mathbf{Y}\mid\mathbf{X})$ we assume that the noise $Z$ is less predictive of $Y$ compared to $E$ if $\mathbf{X}$ is given. This can be roughly interpreted as the noise does not prove useful for predicting $Y$ from $\mathbf{X}$ compared to the ground truth environment label.
@@ Appendix D Experiments: General Remarks
$\mathcal{R}_{\text{II}}=\frac{\mathcal{M}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{M}(f^{E\mid\mathbf{X}})}{\mathcal{M}(f^{E\mid\mathbf{X}})}$ (24)
@@
$\mathcal{R}_{\text{III}}=\frac{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}$ (25)
@@
D.1 Computational complexity
We run all experiments using four Titan X GPUs, with 12GB VRAM each. On this hardware, each experiment requires between two and three days to run to completion. Our code base provides several utilities to reduce the overall memory footprint, allowing reproduction of our experiments on less powerful hardware.
@@ E.1 Data Generation
Simpson’s Paradox [3, 4] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined. There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [59] or an evaluation of the efficacy of different treatments for kidney stones [60].
To replicate this, we create a dataset inspired by an illustration of Simpson’s Paradox on Wikipedia [61]. The dataset consists of a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target. Unless otherwise specified, we generate the data by taking an equal number of samples from each mixture component, defining the environment as a one-hot vector over the mixture components.
@@
n_domains 55 number of mixture components
@@
The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [62].
@@ E.2 Training Details
Now, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 7 for one trained model. Our model captures and utilizes the characteristics of each environment for prediction. In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences. Note that we obtained the best results by considering a class of linear models that aligns with the data-generating process. However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see Section E.3).
@@ E.3 Non-Linear Models
In the experiments in Section 4.2, we considered linear models for our model and the baseline. In the following, we show results for the non-linear model class in Figure 8. We compare predictions of a baseline model and our model on all environments in Figure 9. We see that the extrapolation task fails in some cases as in environment 1. This is due to the mismatch of the considered model class and ground truth model.
@@ Appendix G Experiment 2: Details
To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data. Therefore, we follow the approach proposed in Section 2.5. We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$. Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension. To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see Section 2.5 for details).
@@ Appendix H Experiment 3 and 4: Details
For the BikeSharing dataset we consider a simple feed-forward neural network in all models. For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized. Here, we employ the Clip features proposed in [63]. The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions. Given that Clip features considerably simplify the task, we performed a minimal hyper-parameter search and ensured that the inference model had a similar number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
Table 5: Experiment 4. Performance comparison between our model and the baseline, broken down by target domain. We compare their performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models experience a performance drop in the OOD setting, but our model can detect with strong certainty when this is the case. See Appendix K for more details.
@@
Dataset / Set Size ProDAS / $128$ OfficeHome / 44 PACS / 44
@@ Appendix I Comparison of Permutation-Invariant Architectures
Therefore, we construct the toy dataset as described in Section E.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
@@
Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [64] activation. For the set-based methods, this is followed by the respective pooling. We choose $8$ heads for the attention-based model.
Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [65].
@@
This dataset, taken from the UCI machine learning repository [57], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bike share system.